# LibMoE: A Library for Comprehensive Research Mixture of Experts in Large Language Models

### v2

@@ 5 Analysis
Router margin measures the dominance of the top-ranked expert in a Mixture-of-Experts (MoE) routing decision. It is defined as the difference between the top-1 and top-2 routing scores, with a larger margin indicating a stronger preference for the top-1 expert see Appendix D.4. Figure 8 shows the layer-wise router margin across all MoE methods, revealing a consistent trend of increasing margin with model depth in both language modeling and vision–language tasks. This indicates that expert dominance becomes more pronounced in deeper layers. In particular, the router margin is close to zero in layer 0 for most methods, reflecting early training stages where routing is uncertain and experts receive balanced weights. As depth increases, the rising margin signals a transition from generalist processing toward more specialized expert allocation, with the exception of the final layer in vision–language models, where the pattern deviates.
Notably, SharedE-V2 and SharedE-V3 exhibit the steepest increase in router margin in the language modeling setting, suggesting a more decisive routing strategy and potentially stronger expert specialization. In contrast, this trend does not carry over to VLM sparse upcycling, where their router margins are comparable to other methods. We hypothesize that this discrepancy stems from the sparse upcycling regime. To examine this, we conduct an additional experiment analogous to Section 5a. As shown in Figure 9, when the shared expert is upcycled, the router is relieved from allocating inputs to general-purpose experts and can instead emphasize task-specific experts, resulting in a lower router margin than in models trained from scratch. This behavior is consistent with the routing patterns discussed in Section 5
@@
In sparse upcycling, experts are initially duplicated from dense MLP layers due to low ensemble efficiency during early training. However, do these experts gradually diverge and specialize over time, or do they remain overly similar to each other throughout the training process?. In Figure 10, we investigate the level of similarity between expert output weights at each layer. Higher values indicate greater similarity among experts, suggesting redundancy or representational convergence. The results reveal that most MoE variants exhibit low similarity scores, confirming that while experts may start identically, they gradually diverge and specialize over time during sparse upcycling. Interestingly, we observe that similarity scores tend to be slightly higher at the input and output layers compared to the middle layers. This phenomenon is most pronounced in the MoE++ method, which introduces both copy-experts and zero-experts leading to greater diversity in the intermediate layers. Overall, our analysis indicates that excessive similarity among experts does not persist throughout training, and thus, does not undermine the validity of comparing MoE methods under the sparse upcycling regime.
@@
Figure 12: Expert Co-Activation (ECA) in a Vanilla SMoE (158M) on the language modeling task. Shown are co-activation matrices at Layers 4 and 12 across 10K–100K training steps on the validation set.
h) Expert co-activation over time.
To analyze expert interaction behavior, we visualize the expert co-activation patterns matrix during training. Expert co-activation (ECA) is defined as the proportion of instances in which two specific experts are activated simultaneously, normalized by the total number of activations of one of those experts (Muennighoff et al., 2024). A high ECA value indicates that two specific experts frequently activate together. The formal definition of expert co-activation is defined in Appendix D.5.
As shown in Figure 12, the top co-activated expert pairs in both Layer-4 and Layer-12 of the small (158M) Vanilla SMoE remain relatively stable over the course of training, with only minor variations in pairing strength. This stability suggests that each expert consistently collaborates with a limited subset of other experts, potentially reflecting early-formed cooperative patterns that persist throughout optimization. Similar patterns are observed in other SMoE variants, as detailed in Appendix C.3.
@@ 6 Summary of Key Results
Using LibMoE, we observe that the performance gap among current MoE algorithms is relatively small, as shown in Table2 and Tableă1, suggesting that future MoE designs must deliver more substantial performance improvements to justify their added architectural complexity. However, beyond similar end-task accuracy, our behavioral analysis reveals meaningful differences across algorithms in three key dimensions.
@@ 7 Conclusion
In this work, we presented LibMoE, a unified and extensible framework that enables reproducible research on Mixture-of-Experts models across language and vision–language domains. By integrating seven recent algorithms within standardized pipelines and offering analytical tools for routing, specialization, and load balancing, LibMoE lowers the barrier for experimentation under resource constraints while providing fair and systematic comparisons. Our large-scale empirical study reveals that no single method universally dominates, with performance and stability shaped by task type, initialization, and expert design, underscoring the need for flexible and transparent evaluation. We hope that by bridging the gap between algorithmic innovation and practical deployment, LibMoE will not only accelerate MoE research but also provide the community with a principled foundation for building the next generation of scalable, efficient, and interpretable large models.
@@ A.1 Vision Language Model
LibMoE employs a set of popular benchmarks for vision-language models, including AI2D (Kembhavi et al., 2016), Text VQA Validation (Singh et al., 2019), GQA (Hudson, 2019), Hallusion Benchmark (Guan et al., 2023), MathVista (Lu et al., 2023), MMBenchEN (Liu et al., 2023b), MME  (Fu et al., 2023), MMMU Validation(Yue et al., 2023), MMStar (Chen et al., 2024b), POPE (Li et al., 2023), MME-Real World(Zhang et al., 2024c) and OCR Bench(Liu et al., 2024d). We carefully choose these benchmark to assert the model across several vision-language capabilities such as perception, reasoning, OCR, instruction following, and more. Beyond the performance on standard benchmarks, we analyze the algorithms holistically by evaluating their generalization throughout training, expert selection behaviors, and expert specialization, which we will report in Section 5. For benchmarks requiring GPT-based evaluation, such as MathVista and HallusionBench, we use GPT-4o-mini,version 2024-07-18.
@@
We adopt Phi-3.5 Mini (Abdin et al., 2024) as the language model and SigLIP-SO400M (Zhai et al., 2023) as the vision encoder two widely used backbones in recent multimodal systems. During the VIT stage, we upcycle the dense MLP blocks in both the vision encoder and connector into sparse MoE layers, each composed of $N_{E}{=}6$ experts with top-K=3K{=}3 routing.
@@ A.2 Language Model Pretrain
Dataset. We conduct our experiments on the widely-used SlimPajama dataset (Soboleva et al., 2023), a high-quality corpus curated from RedPajama (Computer, 2023), specifically designed for training large language models (LLMs). SlimPajama has become a standard choice for open LLM research and has been employed in the development of several influential models, including TinyLlama (Zhang et al., 2024b) and BTLM (Cer, 2025), as well as in a range of recent empirical studies (Agarwalla et al., 2024; Gupta et al., 2023). Its adoption across diverse works underscores its value for benchmarking and advancing LLM research.
Tokenizer and Model Architecture. We adopt the SentencePiece tokenizer (Kudo & Richardson, 2018) with byte-pair encoding (BPE), which provides a balance between subword granularity and vocabulary efficiency. generally follow Switch Transformer (Fedus et al., 2022) for our model architecture, adopting a standard Transformer backbone with sparsely activated Mixture-of-Experts (MoE) layers. Each MoE layer replaces the conventional feedforward sublayer and comprises $N$ expert networks (FFNs) and a router mechanism. Table 5 summarizes the comprehensive set of hyperparameters and configurations for both scales and different model variants evaluated in our experiments.
Load Balancing Loss and Router Z-Loss. We adopt the standard load balancing loss from Switch Transformer (Fedus et al., 2022) to penalize imbalances in routing decisions, thereby encouraging more uniform expert utilization. This auxiliary term has been shown to mitigate expert under-utilization and promote balanced load distribution, which improves both convergence and generalization in MoE models. In contrast, we do not incorporate the router z-loss in our experiments (Zoph et al., 2022). Following common practice, we set the weight of the load balancing loss to $\alpha=0.01$. Additional experimental details and design considerations for the auxiliary load balancing loss are provided in Appendix C.4.
Evaluation pipeline. We evaluate our implemented models with the Perplexity score (PPL) and zero-shot performance with nine different downstream tasks: LAMBADA (Paperno et al., 2016), BLiMP (Warstadt et al., 2020), Children’s Book Test (Hill et al., 2015), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-Easy (Clark et al., 2018), RACE (Lai et al., 2017), SIQA (Sap et al., 2019) and CommonSenseQA (Talmor et al., 2018). For LAMBADA, we use the detokenized version from OpenAI, and we evaluate the top-1 accuracy of the last word (it can span multiple tokens; here we use greedy decoding). For CBT, BLiMP, and RACE, we measure the accuracy of each task and report the average accuracy of the tasks.
@@ B.1 Vision-Language Model
To initialize the router network within the MoE blocks, we follow the approach used in the official GPT-2 repository. [^1] https://github.com/openai/gpt-2 Specifically, router weights are drawn from a normal distribution $\mathcal{N}(0,0.02)$ with a fixed random seed of 42 to ensure reproducibility. Only the router is initialized from scratch, while all expert weights are fully trained during VIT. This setup guarantees a controlled and fair evaluation across diverse sparse routing strategies.
@@ B.2 Language Modeling
Scale Model # params # act. params # trained tokens $d_{\mathrm{model}}$ H $d_{\mathrm{head}}$ $N_{E}$ $K_{r}$ $N_{s}$ Expert dim
Small SMoE 158M 36M 6.5B 512 12 82 66 8 0 128
@@
Table 5 summarizes the key hyperparameters, covering model dimensionality, number of attention heads, expert counts, and routing strategies. The small-scale setting processes 6.5B tokens with model dim $d_{model}=512$ and number of attention heads $H=12$, while the large-scale setting extends to 26.2B tokens with $d_{model}=1024$ and $H=16$. Each variant differs in expert number ($N_{E}$), routing capacity ($K_{r}$), and warmup strategies, with expert dimensions set to 128 for small and 256 for large models. These standardized yet diverse configurations ensure a balanced comparison across MoE algorithms while reflecting realistic large-scale training regimes.
Table 6: Training hyperparameter settings for LibMoE across two model scales (158M and 679M parameters) in the language modeling task.
Hyperparameter 158M 679M
@@
Appendix C Addtional Analysis
C.1 Router Saturation
@@
Router Saturation - first introduced in OLMoE (Muennighoff et al., 2024) - quantifies the proportion of overlapping activated experts between an intermediate checkpoint at training step $t$ and the final checkpoint. This metric serves as an indicator of the router’s convergence dynamics throughout training. Higher router saturation values indicate greater alignment in expert selection, signifying that the router’s decisions are becoming increasingly consistent with its final checkpoint. Consequently, router saturation provides insight into the convergence of expert assignments of routing strategies during training. The formal definition and formula are defined in the Appendix D.2.
@@
As shown in Figure 15, router saturation for all evaluated methods rises sharply during training, with most surpassing 60% or even 85% in Hellaswag benchmark saturation within the first 10% of training under top-8 selection. In comparison, top-1 selection exhibits a slightly slower convergence, indicating . Notably, XMoE stands out as an outlier, converging more slowly than other variants, a trend consistent with its relatively lower performance in Table 2. Overall, this early stabilization aligns with previous works (Muennighoff et al., 2024; Xue et al., 2024; Nguyen et al., 2025b; Kang et al., 2025), indicating a general tendency for MoE routers to converge rapidly toward stable expert assignments. Such early convergence is also aligned with the learning strategy advocated by Stable MoE (Dai et al., 2022b), which aims to mitigate the fluctuation problem in expert allocation.
@@
Figure 15: Router Saturation across methods during training on the language modeling task. We present both Top-1 and Top-8 routing results for small (158M) and large (679M) models, illustrating the progression of router convergence across model scales and expert selection strategies.
C.2 Cooperation or Competition
@@
Table 7: Performance variation under different router temperatures across MoE algorithms. We evaluate the impact of router temperature ($\tau$) on the behavior of expert cooperation and competition in both small (158M) and large (679M) models, using two language understanding tasks (BLiMP and HellaSwag).
@@
Small Model (158M) 1.0 BLIMP 77.71% 76.75% 76.53% 77.37% 77.20% 76.91% 77.23%
@@
Large Model (679M) 1.0 BLIMP 80.47% 81.08% 80.38% 80.98% 81.28% 81.21% 80.88%
@@
Table 7 summarizes the effects of varying the router’s temperature parameter on model performance for both small (158M) and large (679M) model sizes across two representative language understanding tasks. We observe that deviating from the original temperature used during training generally results in performance degradation across both small and large model scales. Notably, applying a low temperature ($\tau=0.1$) induces only marginal performance loss, and in several cases - such as with MoE++ and XMoE - yields slight improvements. In contrast, setting a high temperature ($\tau=10.0$) consistently leads to substantial performance drops for all evaluated models, indicating that increased competition among experts impairs the model’s ability to integrate expert knowledge effectively. Interestingly, XMoE demonstrates greater robustness to temperature changes compared to other variants.
@@
C.3 Experts Co-activation overtime
Figure 20, Figure 21, Figure 22, Figure 25, Figure 24, and Figure 23 shows the additional experts co-activation (ECA) of different MoE methods along with SMoE in Section 5h. A consistent behavior is observed in all SMoE methods, where ECA remains relatively stable over the course of training.
C.4 Load Balancing Loss and Router Z-loss Experiment
Figure 16: Benchmark curves during training in language modeling tasks for models with 158M parameters with and without load balancing loss.
Figure 17: Expert selection ratio of SMoE model with and without load balancing loss in language modeling tasks (158M parameters model size).
Table 8: Performance comparisons in the impact of load balancing loss and router z-loss of small model (158M parameters) in language modeling task.
MoE Method PPL $\downarrow$ LAM BADA BLiMP CBT Hella Swag PIQA ARC- Challenge RACE SIQA Common SenseQA Average
SMoE (0.01 lb) 13.63 25.27% 77.71% 84.18% 29.43% 57.94% 21.20% 30.11% 35.62% 24.65% 42.90%
SMoE (no-lb) 13.69 24.90% 76.81% 84.13% 29.38% 57.51% 21.63% 30.26% 35.67% 25.63% 42.88%
SMoE (0.01 lb + 0.001 z-loss) 13.62 25.49% 77.55% 84.23% 29.11% 58.71% 21.89% 29.63% 35.21% 23.91% 42.86%

### v3

@@ 5 Analysis
Notably, SharedE-V2 and SharedE-V3 exhibit the steepest increase in router margin in the language modeling setting, suggesting a more decisive routing strategy and potentially stronger expert specialization. In contrast, this trend does not carry over to VLM sparse upcycling, where their router margins are comparable to other methods. We hypothesize that this discrepancy stems from the sparse upcycling regime. To examine this, we conduct an additional experiment analogous to Section 5a. As shown in Figure 8, when the shared expert is upcycled, the router is relieved from allocating inputs to general-purpose experts and can instead emphasize task-specific experts, resulting in a lower router margin than in models trained from scratch. This behavior is consistent with the routing patterns discussed in Section 5.
@@
In sparse upcycling, experts are initialized by duplicating dense MLP layers, raising the question of whether they remain similar or progressively specialize during training. To examine this, Figure 9 reports the layer-wise cosine similarity of expert output weights, where values close to zero indicate greater divergence between expert outputs. We find that most MoE variants exhibit low similarity, indicating that experts diverge and specialize over time despite identical initialization. Similarity is slightly higher at the input and output layers than in intermediate layers, with the effect most pronounced in MoE++, which employs both copy- and zero-experts and thus encourages greater diversity in middle layers. Overall, excessive expert similarity does not persist, supporting the validity of comparing MoE methods under the sparse upcycling regime.
@@ 6 Summary of Key Results
Using LibMoE, we observe that the performance gap among current MoE algorithms is relatively small, as shown in Table 1 and Table 2, suggesting that future MoE designs must deliver more substantial performance improvements to justify their added architectural complexity. However, beyond similar end-task accuracy, our behavioral analysis reveals meaningful differences across algorithms in three key dimensions.
@@ 7 Conclusion
In this work, we presented LibMoE, a unified and extensible framework that enables reproducible research on Mixture-of-Experts models across language and vision–language domains. By integrating seven recent algorithms within standardized pipelines and offering analytical tools for routing, specialization, and load balancing, LibMoE lowers the barrier for experimentation under resource constraints while providing fair and systematic comparisons. Particularly, all experiments in this study were conducted on a cluster of 4×H100 GPUs, with the longest experiment taking just under 54 hours for VLMs (of which only 35 hours were spent on MoE training) and 43 hours for language model pretraining. Our comprehensive empirical study reveals that no single method universally dominates, with performance and stability shaped by task type, initialization, and expert design, underscoring the need for flexible and transparent evaluation. We hope that by bridging the gap between algorithmic innovation and practical deployment, LibMoE will not only accelerate MoE research but also provide the community with a principled foundation for building the next generation of scalable, efficient, and interpretable large models.
Supplement to “LibMoE: A Library for Comprehensive Research Mixture of Experts in Large Language Models”
@@ A.1 Vision Language Model
LibMoE employs a set of popular benchmarks for vision-language models, including AI2D (Kembhavi et al., 2016), Text VQA Validation (Singh et al., 2019), GQA (Hudson and Manning, 2019), Hallusion Benchmark (Guan et al., 2023), MathVista (Lu et al., 2023), MMBenchEN (Liu et al., 2023b), MME  (Fu et al., 2023), MMMU Validation(Yue et al., 2023), MMStar (Chen et al., 2024b), POPE (Li et al., 2023), MME-Real World(Zhang et al., 2024c) and OCR Bench(Liu et al., 2024d). We carefully choose these benchmark to assert the model across several vision-language capabilities such as perception, reasoning, OCR, instruction following, and more. Beyond the performance on standard benchmarks, we analyze the algorithms holistically by evaluating their generalization throughout training, expert selection behaviors, and expert specialization, which we will report in Section 5. For benchmarks requiring GPT-based evaluation, such as MathVista and HallusionBench, we use GPT-4o-mini,version 2024-07-18.
@@
We adopt Phi-3.5 Mini (Abdin and others, 2024) as the language model and SigLIP-SO400M (Zhai et al., 2023) as the vision encoder two widely used backbones in recent multimodal systems. During the VIT stage, we upcycle the dense MLP blocks in both the vision encoder and connector into sparse MoE layers, each composed of $N_{E}{=}6$ experts with top-K=3K{=}3 routing.
@@ A.2 Language Model Pretrain
Dataset.
We conduct our experiments on the widely-used SlimPajama dataset (Soboleva et al., 2023), a high-quality corpus curated from RedPajama (Computer, 2023), specifically designed for training large language models (LLMs). SlimPajama has become a standard choice for open LLM research and has been employed in the development of several influential models, including TinyLlama (Zhang et al., 2024b) and BTLM (Dey et al., 2023), as well as in a range of recent empirical studies (Agarwalla et al., 2024; Gupta et al., 2023). Its adoption across diverse works underscores its value for benchmarking and advancing LLM research.
Tokenizer and Model Architecture.
We adopt the SentencePiece tokenizer (Kudo and Richardson, 2018) with byte-pair encoding (BPE), which provides a balance between subword granularity and vocabulary efficiency. generally follow Switch Transformer (Fedus et al., 2022) for our model architecture, adopting a standard Transformer backbone with sparsely activated Mixture-of-Experts (MoE) layers. Each MoE layer replaces the conventional feedforward sublayer and comprises $N$ expert networks (FFNs) and a router mechanism. Table 6 summarizes the comprehensive set of hyperparameters and configurations for both scales and different model variants evaluated in our experiments.
Load Balancing Loss and Router Z-Loss.
We adopt the standard load balancing loss from Switch Transformer (Fedus et al., 2022) to penalize imbalances in routing decisions, thereby encouraging more uniform expert utilization. This auxiliary term has been shown to mitigate expert under-utilization and promote balanced load distribution, which improves both convergence and generalization in MoE models. In contrast, we do not incorporate the router z-loss in our experiments (Zoph et al., 2022). Following common practice, we set the weight of the load balancing loss to $\alpha=0.01$. Additional experimental details and design considerations for the auxiliary load balancing loss are provided in Appendix D.4.
Evaluation pipeline.
We evaluate our implemented models with the Perplexity score (PPL) and zero-shot performance with nine different downstream tasks: LAMBADA (Paperno et al., 2016), BLiMP (Warstadt et al., 2020), Children’s Book Test (Hill et al., 2015), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-Easy (Clark et al., 2018), RACE (Lai et al., 2017), SIQA (Sap et al., 2019) and CommonSenseQA (Talmor et al., 2018). For LAMBADA, we use the detokenized version from OpenAI, and we evaluate the top-1 accuracy of the last word (it can span multiple tokens; here we use greedy decoding). For CBT, BLiMP, and RACE, we measure the accuracy of each task and report the average accuracy of the tasks.
@@ B.1 Vision-Language Model
Hyperparameter Settings and Training Stages.
[…]
Table 5: MoE configuration details across different methods, where $N$ denotes the total number of experts, $K$ the number of routed (active) experts per token, and $N_{s}$ the number of shared experts.
MoE Method $\bm{N}$ $\bm{K}$ $\bm{N_{s}}$
SMoE 6 3 0
SharedE 6 2 1
MoE++ 8 3 0
TC-MoE 15 3 0
MoE Architecture Configurations.
Table  LABEL:tab:moe_config_e_vlm summarizes the MoE configurations used across different methods, where $N$ denotes the total number of experts, $K$ the number of routed (active) experts per token, and $N_{s}$ the number of shared experts. For standard routing-based methods, including SMoE, $\sigma$-MoE, and XMoE, we adopt a common configuration with $N=6$ experts and top-$K=3$ routing. For shared-expert architectures, we follow the VLM setting with one shared expert ($N_{s}=1$) and reduce the number of routed experts to $K=2$ to maintain comparable computational cost. For MoE++, we follow Jin et al. (2024) and augment the expert set by introducing zero-computation experts, resulting in a total of $N=8$ experts while keeping $K=3$. For TC-MoE, we follow Yan et al. (2025), where the effective expert pool is expanded via ternary compositions, yielding a total of $2N+K$ experts; in our configuration, this corresponds to $N=15$ with $K=3$.
@@ B.2 Language Modeling
Scale Model # params # act. params # trained tokens $d_{\mathrm{model}}$ H $d_{\mathrm{head}}$ NN $K$ $N_{s}$ Expert dim
Small SMoE 0.15B 36M 6.5B 512 12 82 66 8 0 128
@@
Hyperparameter Settings and Training Stages.
Table 6 summarizes the key hyperparameters, covering model dimensionality, number of attention heads, expert counts, and routing strategies. The small-scale setting processes 6.5B tokens with model dim $d_{model}=512$ and number of attention heads $H=12$, while the large-scale setting extends to 26.2B tokens with $d_{model}=1024$ and $H=16$. Each variant differs in expert number ($N$), routing capacity ($K$), and warmup strategies, with expert dimensions set to 128 for small and 256 for large models. These standardized yet diverse configurations ensure a balanced comparison across MoE algorithms while reflecting realistic large-scale training regimes.
Table 7: Training hyperparameter settings for LibMoE across two model scales (0.15B and 0.68B parameters) in the language modeling task.
Hyperparameter 0.15B 0.68B
@@
MoE Architecture Configurations.
Table 6 also specifies the MoE architecture configurations used in the language modeling pre-training setting, where $N$ denotes the total number of experts, $K$ the number of routed (active) experts per token, and $N_{s}$ the number of shared experts. For standard routing-based methods (SMoE, $\sigma$-MoE, and XMoE), we adopt a common configuration with $N=66$ experts and top-$K=8$ routing. For shared-expert variants (SharedE-V2/V3), we set $N_{s}=2$ shared experts and reduce routed capacity to $K=6$ to maintain comparable compute. For MoE++, following Jin et al. (2024), we augment the expert pool by adding zero-computation experts, resulting in $N=66+8$ while keeping $K=8$. For TC-MoE, following Yan et al. (2025), the effective expert pool is expanded via ternary compositions, yielding $2N+K$ experts; under our configuration, this corresponds to $N=66\times 2+8$ with $K=8$. Across both model scales, we keep the expert dimension fixed (128 for 0.15B and 256 for 0.68B) to ensure a controlled comparison across MoE designs.
Appendix C Comparison Between Dense and SMoE Models
Figure 13: Training benchmark curves comparing Dense and SMoE models during language model pre-training (0.67B parameters).
Table 8: Performance comparison between dense and SMoE models for language model pre-training, evaluated on a small-scale model (0.15B parameters) and a large-scale model (0.68B parameters). PPL denotes perplexity; lower values indicate better performance.
MoE Method PPL $\downarrow$ LAMBADA BLiMP CBT Hella Swag PIQA ARC- Easy RACE SIQA Common SenseQA AVG $\uparrow$
Small Model (0.15B) Dense (36M) 17.04 19.74 73.48 81.03 27.56 55.82 32.09 29.36 35.72 25.14 42.22
SMoE 13.63 25.27 77.71 84.18 29.43 57.94 32.68 30.11 35.62 24.65 44.18
Large Model (0.68B) Dense (131M) 11.51 31.00 77.43 87.45 31.90 61.37 35.14 31.58 36.90 27.03 46.64
SMoE 9.51 37.13 80.47 89.83 37.49 64.36 38.22 33.03 37.41 26.54 49.39
Figure 13 and Table 8 presents a comparative analysis between dense baselines and SMoE models in the language model pre-training setting. Across all reported metrics, SMoE models consistently achieve stronger performance during the early stages of training, indicating faster optimization and more efficient utilization of model capacity compared to dense counterparts.
Appendix D Additional Analysis
D.1 Router Saturation
@@
Router Saturation - first introduced in OLMoE (Muennighoff et al., 2024) - quantifies the proportion of overlapping activated experts between an intermediate checkpoint at training step $t$ and the final checkpoint. This metric serves as an indicator of the router’s convergence dynamics throughout training. Higher router saturation values indicate greater alignment in expert selection, signifying that the router’s decisions are becoming increasingly consistent with its final checkpoint. Consequently, router saturation provides insight into the convergence of expert assignments of routing strategies during training. The formal definition and formula are defined in the Appendix E.4.
@@
As shown in Figure 14, router saturation for all evaluated methods rises sharply during training, with most surpassing 60%, and even reaching over 85% on the HellaSwag benchmark, within the first 10% of training under top-8 selection. In contrast, top-1 selection exhibits slightly slower convergence, indicating a more gradual stabilization of expert assignments. Notably, XMoE stands out as an outlier, converging more slowly than other variants—a trend consistent with its relatively lower performance reported in Table 2. Overall, this early stabilization behavior aligns with prior work (Muennighoff et al., 2024; Xue et al., 2024; Nguyen et al., 2025b; Kang et al., 2025), suggesting a general tendency for MoE routers to converge rapidly toward stable expert assignments. Such early convergence is also consistent with the learning strategy advocated by Stable MoE (Dai et al., 2022b), which aims to mitigate fluctuations in expert allocation during training.
@@
Figure 14: Router Saturation across methods during training on the language modeling task. We present both Top-1 and Top-8 routing results for small (0.15B) and large (0.68B) models, illustrating the progression of router convergence across model scales and expert selection strategies.
D.2 Cooperation or Competition
@@
Table 9: Performance variation under different router temperatures across MoE algorithms. We evaluate the impact of router temperature ($\tau$) on the behavior of expert cooperation and competition in both small (0.15B) and large (0.68B) models, using two language understanding tasks (BLiMP and HellaSwag).
@@
Small Model (0.15B) 1.0 BLIMP 77.71% 76.75% 76.53% 77.37% 77.20% 76.91% 77.23%
@@
Large Model (0.68B) 1.0 BLIMP 80.47% 81.08% 80.38% 80.98% 81.28% 81.21% 80.88%
@@
Table 9 summarizes the effects of varying the router’s temperature parameter on model performance for both small (0.15B) and large (0.68B) model sizes across two representative language understanding tasks. We observe that deviating from the original temperature used during training generally results in performance degradation across both small and large model scales. Notably, applying a low temperature ($\tau=0.1$) induces only marginal performance loss, and in several cases - such as with MoE++ and XMoE - yields slight improvements. In contrast, setting a high temperature ($\tau=10.0$) consistently leads to substantial performance drops for all evaluated models, indicating that increased competition among experts impairs the model’s ability to integrate expert knowledge effectively. Interestingly, XMoE demonstrates greater robustness to temperature changes compared to other variants.
@@
D.3 Expert Co-Activation Over Time
Figure 15: Expert co-activation (ECA) matrices for a Vanilla SMoE (0.15B) on the language modeling task, shown at Layers 4 and 12 across training checkpoints (10K–100K steps) on the validation set.
To further examine expert interaction dynamics, we analyze the evolution of expert co-activation (ECA) patterns throughout training. Following Muennighoff et al. (2024), ECA measures how frequently two experts are activated together, normalized by the total number of activations of one expert (see Appendix E.5 for the formal definition). Higher ECA values indicate persistent co-utilization between expert pairs.
Figure 15 shows the ECA matrices for Layer 4 and Layer 12 of a small (0.15B) Vanilla SMoE model at multiple training checkpoints. Across training, the identity of the most strongly co-activated expert pairs remains largely unchanged, with only modest fluctuations in co-activation strength. This consistency suggests that expert collaboration structures emerge early in training and remain stable thereafter, indicating limited reorganization of expert interactions during later optimization.
We observe similar stability in expert co-activation patterns across other SMoE variants, including XMoE, SMoE-Sigmoid, DeepSeek-V2, DeepSeek-V3, TCMoE, and MoE++. Corresponding ECA visualizations are provided in Figure 20, Figure 21, Figure 22, Figure 25, Figure 24, and Figure 23. Collectively, these results indicate that, despite differences in routing mechanisms and architectural refinements, expert co-activation relationships in SMoE-style models are remarkably stable over training.
D.4 Load Balancing Loss and Router Z-loss Experiment
Figure 16: Benchmark curves during training in language modeling tasks for models with 0.15B parameters with and without load balancing loss.
Figure 17: Expert selection ratio of SMoE model with and without load balancing loss in language modeling tasks (0.15B parameters model size).
Table 10: Performance comparisons in the impact of load balancing loss and router z-loss of small model (0.15B parameters) in language modeling task.
[…]